OpenStudy (anonymous):

One side of a rectangle is 3 inches shorter than the other side, and the perimeter is 54 inches. Which of the following equations could be used to determine the dimensions of the rectangle? 2x + 2(x - 3) = 54 2x + 2x - 3 = 54 2x - 3 = 54

OpenStudy (anonymous):

one side=x other=x-3 perimeter =54 (x-3)*2 +(x)=54 this is equal to first one so thats the answer 2x-6+2x=54 4x=60 x=15

OpenStudy (anonymous):

(x-3)*2 + 2*(x)=54 missed the 2 *

OpenStudy (anonymous):

so the second or first one?

OpenStudy (anonymous):

first.

OpenStudy (anonymous):

thanks